The mass of the empty container is 12.70g and when filled with gas, was 18.20g. The pressure inside the container is 0.80 atm, temperature 25.40 degrees and volume is 5.40L

The moles (mol) of gas=
The molecular mass (g/mol) of the gas=
1
Expert's answer
2020-11-20T08:09:04-0500

Pressure inside the container, P = 0.80 atm

Temperature = 25.40°C

So, T = (273+25.40)K

= 298.40 K

Volume of container, V = 5.40 L


We know Ideal gas equation,

PV = nRT

Where, R = universal gas constant

= 0.082 L.atm.mol–1.K–1


From above ideal gas equation,

Number of moles of gas,

n = PV/RT

Or, n = (0.80×5.40)/(0.082×298.40)

= 0.281 mole

Hence, The moles (mol) of gas = 0.281 mole


Let, The molar mass of gas = M g/mol



The mass of container filled with gas

= 8.20g.

The mass of the empty container

= 12.70g

So, The mass of gas,

W = (18.20 – 12.70)g

= 5.50 g

Number of moles of gas,

n = (W/M) mole

= (5.50/M) mole

Now, n = 0.281 mole

So, (5.50/M) = 0.281

Or, M = 5.50/0.281

Or, M = 19.573 g/mol


Hence, The molecular mass (g/mol) of the gas = 19.573 g/mol